# How precise molecular dynamic simulations are and what we can learn from it?

### 21 Dec 2023 v1

@@ Abstract
We investigated the influence of a finite number of computer simulations by molecular dynamics method on fluctuations in thermodynamic properties. As a case study, we used the two-dimensional Lennard-Jones system. Besides being an archetypal system, the 2D Lennard-Jones system is a subject of long debate, as to whether its melting transition is continuous (infinite-order) or discontinious ( first-order) one. We have discovered that anomalies on the equation of state (the van-der-Waals or Myer-Wood loops) previously considered a hallmark of first order phase transition, are at best at the level of noise, since they have the same magnitude as the amplitude of pressure fluctuations. Having calculated the dependence of temperature fluctuations on the number of particles, we obtained another strong evidence of continuous character of this melting transition, which manifests itself in continuous change of heat capacity across the melting curve. We also compared the results obtained for temperature fluctuations with prediction made by statistical mechanics .
Fluctuations are an inevitable feature of finite size systems’ thermal behavior, which manifested itself in the Brownian motion of small particles. In thermodynamic limit (infinite size of the system), fluctuations decrease, so we can safely deal with the system’s average thermodynamic parameters.
However, when we apply molecular dynamics and in more general sense, any computer simulations, we use quite small finite size systems. Inevitable question arises how close we are to the thermodynamic limit. The answer to this question is an estimation of system’s fluctuations.
This problem previously was acknowledged in the paper of Hickman and Mishin Hickman and Mishin (2016), but they considered only the temperature fluctuations. Their description starts from outline of fundamental problem with temperature fluctuations in computer simulations which is lacking unanimously adopted explanation. This explanation can be found in the well-known course of theoretical physics Landau et al. (1980) which produces a result that relative fluctuations of thermodynamic parameter are roughly inversely proportional to the square root of the number of particles in the system.
If we restrict ourselves to description of NVE ensemble only (with constant number of particles N, volume V and energy E ), the pressure and temperature of the system can be expected to fluctuate around some average values. Classical statistical mechanics predicts Landau et al. (1980) that fluctuations of pressure P and temperature T of large enough system are distributed according to the Gauss law with standard deviations:
@@
$\Delta P^{2}=k_{B}T\left(\frac{\partial P}{\partial V}\right)_{S}$ (2)
@@
Here, $k_{B}$ – the Boltzmann constant, $c_{V}$ – specific heat at constant volume per particle, index S means constant entropy. The second formula in the case of ideal gas gives:
$\left(\frac{\Delta P}{P}\right)^{2}=\frac{\gamma}{N}$ (3)
@@
where $\gamma$ is an adiabatic exponent. This suggests the general law that relative fluctuations of thermodynamic parameters approximately scale to the square root of the system size. So for typical system size used in molecular dynamics simulations ($N\leq 10^{5}$), relative fluctuations are slightly less than 1 %.
@@
Figure 1: Fluctuations of temperature (upper panel) and relative pressure (lower panel) at $T=0.9$ and two different densities $\rho=1.23$ (blue curves) and $\rho=0.86$ (red curves) which correspond to deep inside crystal phase and just below transition in the liquid phase respectively.
@@
Figure 2: Distribution density of thermal fluctuations at $T=0.9$ and different densities with $N=10^{4}$ particles calculated in NVE ensemble. Density values encoded in the color key at the right side of the plot
@@
Figure 3: Equation of state at different temperatures in the vicinity of melting transition. Inset shows blow up of transition at $T=0.9$ calculated in different ensembles (implicitly assumed NVE ensemble) and various numbers of particles $N$ (see list in the main panel).
@@
To test these assumptions we have chosen the two dimensional Lennard-Jones (2DLJ) system. Beside its archetypical character in computer simulations, its type of crystal melting attracts much attention . Original general theory of 2D crystal described in the seventieth in the works later contributed to the Berezinskii-Kosterlitz-Thouless-Halperin-Nelson-Young theory suggests that melting is infinite-order, two-stage transition between crystal and liquid (the first stage – crystal to hexatic phase transition with the loss of translational order, and the second stage – hexatic to liquid, with the loss of orientation order, see e.g. Ryzhov et al. (2023)). For simplicity we will call it continuous model. However, almost immediately this model was challenged by computer simulation study Barker et al. (1981) suggesting first order phase transition in 2DLJ system. Subsequently this system was addressed in the number of works Frenkel and McTague (1979); Toxvaerd (1981); Tobochnik and Chester (1982); Koch and Abraham (1983); Bakker et al. (1984); Udink and van der Elsken (1987); Chen et al. (1995); Somer et al. (1997, 1998) which produced controversial opinions ranging from first-order to continuous character of melting transition. Hot debates followed through 1980-1990s but did not lead to a clear conclusion. Echoes of the debates continue until recentlyWierschem and Manousakis (2010); Patashinski et al. (2010); Wierschem and Manousakis (2011); Tsiok et al. (2022). Still, it is worth mentioning that first order was predicted by the authors who used rather a small system $N<10^{4}$ particles, continuous transition was claimed by the authors who used a larger system up to $N=10^{5}$ and intermediate position was taken by authors for the system with $N=2500$ particles Patashinski et al. (2010). The most recent paper using $N=512^{2}$ particles Tsiok et al. (2022) was in favor of first-order transition. It seems that authors’ decisions are unconsciously influenced by scale of fluctuations . Most important of them are pressure fluctuations which influence the shape of equation of state (P-V) curve. Usually the presence of van-der-Waals loops on the P-V curve (or equivalently the Mayer-Wood loop on $P-\rho$ where $\rho$ is numerical density ) is considered a decisive indicator of the first order transition. Still, the depth of the loops reported in many of the works cited, is rather shallow and does not allow to make a decision in favor of first-order or continuous phase transition . The most convincing would be comparison of the effect observed with the inherent system noise due to thermal fluctuations.
@@
We tested validity of theoretical predictions by computations with classical molecular dynamics simulations of 2DLJ system as implemented in LAMMPS software package Plimpton (1995); Plimpton et al. . Input files from Carsten Svaneborg group home page cs: were adapted to our needs. We used NVE ensemble with Langevin thermal dynamics with time step 0.01 LJ units. We used $510^{4}$ steps for thermal equilibration and next $510^{4}$ steps for estimation of thermodynamic parameters and their fluctuations . Validity of this choice will be explained later. As a rule we used $N=10^{4}$ particles. Some of the results were checked with larger system containing $N=10^{5}$ particles. Also longer simulation times and NVT ensemble were used. Initial configuration of particles was purely random. Changing the ensemble to NVT with Nose-Hoover thermostat slightly influenced the result making the system fluctuate around initially set temperature but the amplitude of fluctuations did not change. Starting configuration of particles was chosen random with ititial minimization of energy to avoid overlapping particles. The temperature varied in the range $T=0.6-1.1$ with step 0.1 and density range $\rho=0.55-1.25$. To compare our findings with previous results the temperature $T=3.0$ near the melting transition was investigated . The distribution density of thermal fluctuations was estimated by density function with default parameters as implemented in R software package R Core Team (2012).
@@
Figure 7: Equation of state in vicinity of melting temperature calculated in the current work for NVE ensemble with $N=10^{5}$ particles at $T=3.0$ (green dashed curve is guide to the eyes). Blue curve is the data for NVT ensemble for $N=256^{2}$ particles at the same temperature obtained in Ref. Tsiok et al. (2022). Inset shows fluctuations of temperature for the chosen (marked in red) point in the middle of the melting transition.
Figure 8: Heat capacity of the 2DLJ system estimated from thermal fluctuations. Dashed curve is guide to the eyes.
Fluctuations of pressure and temperature of system with $N=10^{4}$ particles at T=0.9 are shown in Fig. 1 at the same set temperature $T\approx 0.9$ and different densities ($\rho=1.23$ and $\rho=0.86$) which correspond to part of the transition deep inside the crystal phase and the liquid part just below the transition. As it was expected, fluctuations of temperature reach about 1% with almost normal distribution (see Fig. 2). The P-$\rho$ curves in vicinity of these density-temperature conditions (near the melting transition) are shown in Fig. 3. It is clear, that no van-der-Waals loops are observed in this region and with precision of pressure fluctuations might be considered as monotonous function increasing with the density rise. Still we need to inspect fluctuations of temperature to make decisive conclusions.
@@
For comparison we provide distributions densities obtained for the same number of particles in the same conditions but calculated in the NVT ensemble. As expected for this sort of ensemble the distribution maximum (mean value) is closer to the initially set temperature value. However, the standard deviation is almost the same as in NVE ensemble. Beside that is interesting to note that at $\rho\approx 1$ the distribution is also bimodal. This fact possibly catches an important feature in the dynamic of the system presumably arising from transition from the crystal to hexatic phase.
@@
To demonstrate this idea we provide the results for the larger NVE ensemble with $N=10^{5}$ particles at the same temperature $T=0.9$ in the nearest vicinity of the transition (see Fig. 5). As expected, with the increase of the particles number, the values of temperature and pressure fluctuations are diminished but the slow relaxation of pressure is conserved in the crystal part of transition. However, at the upper half of time range investigated the relaxation almost reaches its equilibrium value which is demonstrated by the density distribution of temperature fluctuations in this time region (see Fig. 6). It is clear that mean temperature is only slightly above the initially set one and the drift of the mean temperature is negligible in comparison to the standard deviation of fluctuations. This means that the absence of the van-der-Waals loops on the pressure-density curve observed in Fig. 5 is not related to the drift of temperature at higher densities.
@@
To compare our findings with previous results  Tsiok et al. (2022) we performed calculation at higher temperature T=3.0 in NVE ensemble and $N=10^{5}$ particles (see Fig. 7). In the work of Tsiok et al. Tsiok et al. (2022) calculation was done in NVT ensemble at the same temperature, but with larger particle number of $N=256^{2}$. It was stated there that the van-der-Waals loops were observed, which demonstrate first-order phase transition at least at one stage of transition (hexatic-liquid one). Still, the van-der-Waals loop observed there had rather small amplitude (in Fig. 7 it looks like a plateau), being significantly smaller than the standard deviations of pressure fluctuations calculated by us. So, there is not enough evidence of a first order of hexatic-liquid transition to overthrow existing theory of 2D melting. By the way fluctuations observed in the middle of transition, shown in the inset of Fig. 7 demonstrate convergence of pressure relaxation at longer time range ($t=210^{5}$). From this demonstration it is evident that the time range used by us (upper half of $t=10^{5}$ range) is large enough to fully converge pressure values in vicinity of the melting transition.
@@
There is qualitative difference between the two curves in Fig. 7 – the plateau on the curve of Tsiok et al. Tsiok et al. (2022) is not only flatter than the one reported in present work, but also spans a longer range of densities $\rho$. To explain this discrepancy, we can refer to the difference between initial conditions in our work and Ref. Tsiok et al. (2022). In our setup initial atom coordinates were purely random but in Ref. Tsiok et al. (2022) they started from ideal triangular lattice. The latter case favors stability of the crystal phase, so it persists longer below melting transitions which results in van-der-Waals loops on equation of state curve. Flatness of this curve depends on the number of particles in the system so at the thermodynamic limit (the infinite number of particles) the van-der-Waals loop is believed to transform into ideal plateau with zero derivative in the middle of transition. It was demonstrated for the first time for the 2DLJ system Toxvaerd (1981) that deepness of the van-der-Waals loop in the system with “crystal” initial state is correlated to the number of particles – the larger the number the flatter curve. From this work it can be learned that the amplitude of van-der-Waals loop oscillation has the same order of magnitude as system fluctuations . In Ref. Toxvaerd (1981) for $N=256$ system both values are about 10%, which corresponds to the general rule of thumb that relative fluctuations are inversely proportional to the square root of particles number.
@@
There is another argument in favor of first-order phase transition, namely phase coexistense observed in computer experiments between hexatic and liquid phase in the region of “van-der-Waals loop” Tsiok et al. (2022). We did not investigate this problem closely but previous report obtained by Patashinskii et al. Patashinski et al. (2010) in $N=2500$ 2DLJ system demonstrates that this separation has transient character, so the regions of hexatic phase and translational disordered liquid freely transform between each other. In our opinion (we agree with authors of Ref. Patashinski et al. (2010)) taht such a transformation is an indication of continuous transition between the two phases.

### 02 Mar 2024 v2

# The precision of molecular dynamics simulations and what we can learn from it?

@@ Abstract
We have investigated by molecular dynamics method the influence of a finite number of particles used in computer simulations on fluctuations of thermodynamic properties. As a case study, we used the two-dimensional Lennard-Jones system. 2D Lennard-Jones system, besides being an archetypal one, is a subject of long debate, as to whether it has continuous (infinite-order) or discontinuous (first-order) melting transition. We have found that anomalies on the equation of state (the van-der-Waals or Myer-Wood loops) previously considered a hallmark of the first order phase transition, are at best at the level of noise, since they have the same magnitude as the amplitude of pressure fluctuations. So, they could be regarded as statistically unsignificant effect. Also, we estimated inherent statistical noise, present in computer simulations, and came to conclusion, that it is larger, than predicted by statistical physics, and the difference between them (called algorithmic fluctuations) is possibly due to the computer-related issues. It was demonstrated that these fluctuations in principle could be observed in real-life physical experiments which would lead to practical resolution of The Matrix hypothesis.
I Introduction
Fluctuations are an inevitable feature of finite size systems’ thermal behavior, which manifested itself in the Brownian motion of small particles. In thermodynamic limit (infinite size of the system), fluctuations decrease, so we can safely deal with average thermodynamic parameters.
However, molecular dynamics and, in general, any computer simulations, use quite small finite size systems. Inevitable question arises with this, how close we are to the thermodynamic limit . To answer this question we have to estimate the system’s fluctuations.
The problem was previously acknowledged in Hickman and Mishin paperHickman and Mishin (2016), but they considered the temperature fluctuations only. Their description starts from outline of fundamental problem with temperature fluctuations in computer simulations, which has no unanimous explanation. The explanation can be found in well-known course of theoretical physics Landau et al. (1980) which gives a result that relative fluctuations of thermodynamic parameter are roughly inversely proportional to the square root of the number of system’s particles.
If we limit ourselves to description of NVE ensemble (with constant number of particles N, volume V and energy E ), the pressure and temperature of the system are expected to fluctuate around some average values. Classical statistical mechanics predicts Landau et al. (1980) that fluctuations of pressure P and temperature T of large enough system are distributed according to the Gauss law with standard deviations:
@@ I Introduction
$\Delta P^{2}=-k_{B}T\left(\frac{\partial P}{\partial V}\right)_{S}=\frac{k_{B}T}{V}B_{S}$ (2)
@@
Here, $k_{B}$ – the Boltzmann constant, $c_{V}$ – specific heat at constant volume per particle, $B_{S}$ – adiabatic bulk modulus, index S means constant entropy. The second formula in the case of ideal gas gives:
$\left(\frac{\Delta P}{P}\right)^{2}=\frac{\gamma}{N}$
@@
where $\gamma$ is an adiabatic exponent. This suggests the general law that relative fluctuations of thermodynamic parameters approximately scale to the square root of the system size. This conclusion is corroborated by computer simulations of pressure fluctuations in non-ideal plasmaBystryi et al. (2014). So for typical system size used in molecular dynamics simulations ($N\leq 10^{5}$), relative fluctuations are slightly less than 1 %. The situation gets worse for ab-initio molecular dynamic simulations, where number of atoms doesn’t exceed several hundreds, where fluctuations can reach of about 10 %.
@@
Figure 1: Fluctuations of temperature (upper panel) and relative pressure (lower panel) at $T=0.9$ and two different densities $\rho=1.23$ (blue curves) and $\rho=0.86$ (red curves) which correspond to the states deep inside the crystal phase and in the liquid phase just below the transition respectively.
@@
Figure 2: Distribution density of thermal fluctuations at $T=0.9$ and different densities with $N=10^{4}$ particles calculated in NVE ensemble. Density values are encoded in the color key at the right side of the plot
@@
Figure 3: Equation of state at different temperatures in the vicinity of melting transition. The inset shows enlarged region in the vicinity of transition at $T=0.9$ calculated for different ensembles and numbers of particles $N$ (see list in the main panel). The type of ensemble if not indicated in the plot means NVE ensemble.
@@
To test these assumptions we have chosen the two dimensional Lennard-Jones (2DLJ) system lj: . Being an archetypical model in computer simulations, its melting transition type attracts much attention . Original theory of 2D crystal melting dating back to the 70s (the Berezinskii-Kosterlitz-Thouless-Halperin-Nelson-Young theory) suggests that melting is an infinite-order, two-stage transition between crystal and liquid. The first stage – crystal to hexatic phase transition with the loss of translational order, and the second stage – hexatic to liquid, with the loss of orientation order, see e.g. Ryzhov et al. (2023). For this theory Kosterlitz and Thouless in 2016 were awarded with Nobel prize. For simplicity, we will call it a continuous model. However, almost immediately this model was challenged by computer simulation study Barker et al. (1981) suggesting the first order phase transition in 2DLJ system. Subsequently, this system was addressed in the number of works Frenkel and McTague (1979); Toxvaerd (1981); Tobochnik and Chester (1982); Koch and Abraham (1983); Bakker et al. (1984); Udink and van der Elsken (1987); Chen et al. (1995); Somer et al. (1997, 1998) which produced controversial opinions ranging from first-order to continuous character of melting transition. Hot debates followed through 1980-1990s but did not lead to a clear conclusion. Echoes of the debates continue until recentlyWierschem and Manousakis (2010); Patashinski et al. (2010); Wierschem and Manousakis (2011); Hajibabaei and Kim (2019); Li and Ciamarra (2020); Tsiok et al. (2022). Still, it is worth mentioning that first order was predicted by the authors who used rather a small system $N<10^{4}$ particles, continuous transition was claimed by the authors who used a larger system up to $N=10^{5}$ and intermediate position - for the system with $N=2500$ particles Patashinski et al. (2010). The most recent paper using $N=512^{2}$ particles Tsiok et al. (2022) favored the first-order transition. It seems that the authors’ opinions are unconsciously effected by the scale of fluctuations . Most important are pressure fluctuations which influence the shape of equation of state (P-V) curve. Usually presence of van-der-Waals loops on the P-V curve (or the Mayer-Wood loop on $P-\rho$ where $\rho$ is numerical density ) is considered a decisive indicator of the first order transition. Still, the depth of the loops reported in many of these works, is rather shallow and does not allow to choose between the first-order or continuous phase transition . The most convincing would be comparison of the effect observed with the inherent system noise due to thermal fluctuations.
II Methods
@@
We tested validity of theoretical predictions by computations with classical molecular dynamics simulations of 2DLJ system as implemented in state of art LAMMPS software package Plimpton (1995); Plimpton et al. . Input files from Carsten Svaneborg group home page cs: were adapted to our needs. We used NVE ensemble with Langevin thermal dynamics with time step 0.01 LJ units. We used $5\cdot 10^{4}$ steps for thermal equilibration and next $5\cdot 10^{4}$ steps for estimation of thermodynamic parameters and their fluctuations . Validity of this choice will be explained later. As a rule we used $N=10^{4}$ particles. Some of the results were checked with larger system containing $N=10^{5}$ particles. Also longer simulation times and NVT ensemble were used. Changing the ensemble to NVT with Nose-Hoover thermostat slightly influenced the result making the system fluctuate around initially set temperature, but the amplitude of fluctuations did not change. Starting configuration of particles was chosen random with initial minimization of energy to avoid overlapping particles. Random configuration taken by us illustrates crystallization of the liquid rather than crystal melting. The order of crystal-liquid phase transition should not be influenced by the direction of the phase transition. The temperature was varied in the range $T=0.6-1.1$ with step 0.1 and density range $\rho=0.55-1.25$. To compare our findings with previous results the temperature $T=3.0$ near the melting transition was investigated . The distribution density of thermal fluctuations was estimated by density function with default parameters as implemented in R software package R Core Team (2012).
III Results
@@
Figure 7: Equation of state in vicinity of melting temperature calculated in the current work for NVE ensemble with $N=10^{5}$ particles at $T=3.0$ (green dashed curve is guide to the eyes). Blue curve is the data for NVT ensemble for $N=256^{2}$ particles at the same temperature obtained in Ref. Tsiok et al. (2022). Inset shows temperature fluctuations for the chosen (marked in red) point in the middle of the melting transition.
Figure 8: Heat capacity of the 2DLJ system estimated from thermal fluctuations by Eq. 1. Dashed curve is guide to the eyes.
Figure 9: Adiabatic bulk modulus $B_{S}$ and heat capacity $C_{V}$ evaluated from Eqs. 1-3. “True” $C_{V}$ and isothermal bulk modulus $B_{T}$ are evaluated by numerical differentiation of internal energy along isochores and equation of state $P(\rho)$ along isotherm $T=0.9$ respectively.
Figure 10: The pressure relaxation curves (left panel) and equation of state at $T=0.9$ (right panel) in three dimensional Lennard-Jones system. Dashed red curves at the right panel are guides to the eyes.
Figure 11: Adiabatic bulk modulus $B_{S}$ and heat capacity $C_{V}$ evaluated from Eqs. 1-3 in 3D case. Isothermal bulk modulus $B_{T}$ are evaluated by numerical differentiation of equation of state $P(\rho)$ along isotherm $T=0.9$.
Fluctuations of pressure and temperature of system with $N=10^{4}$ particles at T=0.9 are shown in Fig. 1 at the same set temperature $T\approx 0.9$ and different densities ($\rho=1.23$ and $\rho=0.86$) which correspond to thermodynamic conditions deep inside the crystal phase and the liquid phase just below the transition. As it was expected, fluctuations of temperature reach about 1% with almost normal distribution (see Fig. 2). The P-$\rho$ curves in vicinity of these density-temperature conditions (near the melting transition) are shown in Fig. 3. It is clear, that no van-der-Waals loops are observed in this region and with precision of pressure fluctuations they might be considered as monotonous functions increasing with the density rise. Still we need to inspect fluctuations of temperature to make decisive conclusions.
@@
For comparison we provide distributions densities obtained for the same number of particles in the same conditions but calculated in the NVT ensemble (Fig. 4). As expected for this sort of ensemble the distribution maximum (mean value) is closer to the initially set temperature value. However, the standard deviation is almost the same as in NVE ensemble. It is interesting to note that at $\rho\approx 1$ the distribution is also bimodal. This fact possibly catches an important feature in the dynamic of the system presumably arising from transition from the crystal to hexatic phase.
@@
To demonstrate this idea we provide the results for the larger NVE ensemble with $N=10^{5}$ particles at the same temperature $T=0.9$ in the nearest vicinity of the transition (see Fig. 5). As expected, with the increase of the particles number, the values of temperature and pressure fluctuations diminish, but slow relaxation of pressure is conserved in the crystal side of transition. However, at the upper half of time frame used for calculation of thermodynamic parameters the relaxation almost reaches its equilibrium value which is demonstrated by the density distribution of temperature fluctuations in this time region (see Fig. 6). It is clear that mean temperature is only slightly above the initially set one and the drift of the mean temperature is negligible in comparison to the standard deviation of fluctuations. This means that the absence of the van-der-Waals loops on the pressure-density curve observed in Fig. 5 is not related to the drift of temperature at higher densities.
@@
To compare our findings with previous results Tsiok et al. (2022) we performed calculation at higher temperature T=3.0 in NVE ensemble and $N=10^{5}$ particles (see Fig. 7). In the work of Tsiok et al. Tsiok et al. (2022) calculation was done in NVT ensemble at the same $T=3.0$ temperature, but with larger particle number of $N=256^{2}$. It was stated there that the van-der-Waals loops were observed, which demonstrate first-order at least at hexatic-liquid stage of the transition. Still, the van-der-Waals loop observed there had rather small amplitude (in Fig. 7 it looks like a plateau), being significantly smaller than the standard deviations of pressure fluctuations calculated by us.
By the way, fluctuations observed in the middle of transition, shown in the inset of Fig. 7 demonstrate convergence of pressure relaxation at longer time range ($t=2\cdot 10^{5}$). From this demonstration it is evident that the time range used by us (upper half of $t=10^{5}$ interval) is large enough to fully converge the pressure values in the vicinity of melting transition.
@@
There is qualitative difference between the two curves in Fig. 7 – plateau on the curve of Tsiok et al. Tsiok et al. (2022) is not only flatter than the one reported in present work, but also spans a longer range of densities $\rho$. To explain this discrepancy, we can refer to the difference between initial conditions in our work and Ref. Tsiok et al. (2022). In our setup initial atom coordinates were purely random but in Ref. Tsiok et al. (2022) they started from ideal triangular lattice. The latter case favors stability of crystal phase, so it persists longer below melting transitions which results in van-der-Waals loops on equation of state curve. Flatness of this curve depends on the number of system’s particles, so at the thermodynamic limit (the infinite number of particles) the van-der-Waals loop is believed to transform into ideal plateau with zero derivative in the middle. It was demonstrated for the first time for the 2DLJ system Toxvaerd (1981) that deepness of the van-der-Waals loop in the system with “crystal” initial state is correlated to the number of particles – the larger the number, the flatter the curve. From this work one can see that the amplitude of van-der-Waals loop has the same order of magnitude as system fluctuations . In Ref. Toxvaerd (1981) for $N=256$ system both values are about 10%, which corresponds to the rule of thumb that relative fluctuations are inversely proportional to square root of the number of particles .
@@
There is another argument in favor of the first-order phase transition, namely phase coexistence observed in computer experiments between hexatic and liquid phase in the region of “van-der-Waals loop” Tsiok et al. (2022). We did not investigate this problem closely, but previous report obtained by Patashinskii et al. Patashinski et al. (2010) in $N=2500$ 2DLJ system demonstrates that this separation has transient character, so that regions of hexatic phase and translational disordered liquid freely transform to each other. In our opinion (we agree with authors of Ref. Patashinski et al. (2010)) such a transformation is an indication of continuous transition between the two phases.
There were two recent papers proposing different scenarios of the first order phase transition Hajibabaei and Kim (2019); Li and Ciamarra (2020). Rather than discussing contradictions in various simulations we would like to draw attention to the point – how many sigmas the effect reported (deepness of the van-der-Waals loops) is? In other words, we insist on the importance of being earnest with the error bars. Comparison with the noise band of pressure relaxation curve reported in our paper demonstrates that it is also much below 1-sigma. So, there is no point to discuss such a negligible effect. The effect should be as much as 3-sigma, to be statistically significant. However, effect and noise in 2DLJ system are correlated – the larger the system size , the smaller the noise, and the shallower van der Waals loops Toxvaerd (1981). In fact, this is discussed in Hajibabaei and Kim (2019) as size effect– the energetic barrier between hexatic phase and liquid diminishes as $1/\sqrt{N}$ with the system size. Hence, in the thermodynamic limit, this means that energetic barrier between two phases vanishes, that is transition is continuous, therefore it should be in full agreement with conclusions of the Berezinskii-Kosterlitz-Thouless-Halperin-Nelson-Young theory. Therefore, there is not enough evidence of a first order of hexatic-liquid transition to overthrow existing theory of 2D melting.
Later we demonstrate that the noise observed in the 2DLJ system at large extent corresponds to the predictions of statistical physics.
IV Discussion
Previously we have demonstrated the finite precision of molecular dynamic simulations. In this section we answer the second part of the question posed in the title – what we can learn from finite precision of molecular dynamics experiments. As it follows from Eqs. 1-2 from fluctuations of temperature and pressure we could estimate isochoric heat capacity and adiabatic bulk modulus.